# Surprisal from Larger Transformer-based Language Models Predicts fMRI Data More Poorly

### v1

@@ 3.2 Predictors
3.3 Regression Modeling
Once the surprisal estimates were collected, we fit a set of linear mixed-effect (LME) models to BOLD signals with lme4 (Bates et al., 2015) using the fit partition. During regression modeling, in addition to surprisal estimates, we also included a set of baseline predictors for both corpora. For Natural Stories fMRI, the baseline predictors were the index of the word position in the sentence and the word length in characters. For Pereira fMRI, the baseline predictors were sentence length in words, index of sentence position within each passage, and index of the passage containing the sentence.
All models were fit using raw BOLD values along with the above baseline predictors. The predictors were centered and scaled before being used to fit the LME models. The following formula used for fitting the LME models for Natural Stories fMRI contains by-subject random slopes for word position (sentpos), word length (wlen), surprisal estimate (surp) of the data point, and a by-subject intercept. An additional random intercept was included for each subject.
BOLD ~ z.(sentpos) + z.(wlen) + z.(surp) +
       (1 + z.(sentpos) + z.(wlen) +
       z.(surp) | subject) + (1 | subject)
For Pereira, we fit the LME models with the following formula:
BOLD ~ z.(sentlen) + z.(passpos) + z.(lastwordsurp) +
       (1 + z.(sentlen) + z.(passpos) +
       z.(lastwordsurp) | subject) + (1 | sentid)
The LME models included by-subject random slopes for sentence length (sentlen), sentence position within passage (passpos), last-word surprisal (lastwordsurp), and a by-subject intercept. These models also included an additional random intercept for each sentence (sentid). We show the predictive power with the model log likelihood calculated on the held-out partition for both datasets.
3.4 Results
Figure 1 presents the relationship between the regression model log likelihood and log-transformed parameter count. The results of Natural Stories fMRI do not show a significant trend in either direction, replicating Oh et al. (2022), but the results on Pereira fMRI show that as parameter count (model size) increases, the regression model log likelihood decreases significantly ($p$ < 0.05 by a permutation test). [^2] Although the results of Natural Stories fMRI on the held-out partition do not show a clear relationship between predictive power and parameter count, the results of it on the exploratory partition do show a significant inverse scaling ($p$ 0.001). This suggests that for brain imaging data, the predictive power of surprisal estimates decreases as the model size increases, which is in line with previous finding on reading times (Oh and Schuler, 2023; Oh et al., 2024).
Figure 1: Predictive power of surprisal estimates from pretrained LMs on BOLD signals of Natural Stories fMRI and Pereira fMRI.
4 Discussion and Conclusion
Recent studies (Oh et al., 2022; Oh and Schuler, 2023; Oh et al., 2024) have observed a robust negative relationship between Transformer-based LMs’ size and the predictive power of surprisal estimates from those LMs on latency-based measures (i.e., self-paced reading times and eye-gaze durations). This work generalizes this result by examining the predictive power of surprisal estimates from Transformer-based models on brain imaging data. The predictive power of surprisal estimates from 17 Transformer-based LMs across three LM families was evaluated on two fMRI datasets. Results show that the inverse scaling between model size and model fit still obtains on brain imaging data, suggesting that this trend is not specific to latency-based measures and can be generalized to neural measures.
Limitations
This study attempts to examine the predictive power of surprisal estimates from Trasnformer-based models on fMRI data. LLMs evaluated in this work are trained on English text and the datasets evaluated are collected from English speakers. Therefore, these findings may or may not be replicated in other languages.
Ethics Statement
The fMRI datasets used in this study are from published work (Shain et al., 2020; Pereira et al., 2018). Details regarding the data collection and relevant procedures are described in respective publications. Since this work focuses on evaluating the predictive power of LLMs on fMRI data, the potential risks and harmful impacts posed by this study on the society appear to be minimal.

### v2

@@ Appendix A Hyperparameters of Models Examined in this Study
Appendix B Maximal LME models
During LME model fitting, we followed Barr et al. (2013) and started with the maximal LME models. We eliminated one random effect at a time until the LME models converge. To ensure that the regression includes both by-subject and by-item random effects, we removed random effects iteratively from the by-subject and by-item effects. That is, if the LME model does not converge, we first removed a random effect from the by-item group; if the LME model still does not converge, we then move on to removing a random effect from the by-subject group, and so on.
Following is the original design of the LME model for Natural Stories fMRI, in which the eliminated random effects are bolded.
BOLD ˜ z.(sentpos) + z.(wlen) + z.(surp) + (1 + z.(sentpos) + z.(wlen) + z.(surp) | subject) + (1 + z.(sentpos) + z.(wlen) + z.(surp) | storyid)
As for Pereira fMRI, the original design of the LME model is the following, with the eliminated random effect bolded:
BOLD ˜ z.(sentlen) + z.(passpos)+ z.(lastwordsurp) + (1 + z.(sentlen) + z.(passpos) + z.(lastwordsurp) | subject) + (1 + z.(sentlen) + z.(passpos) + z.(lastwordsurp) | sentid)